# Hydrogen reaction rate modeling based on convolutional neural network for large eddy simulation

### v1

@@ 2.3.2 Filtering and downsampling
A spatial filtering is performed to obtain $\overline{\rho}$, $\overline{\rho Y_{\mathrm{H_{2}}}}$, $\overline{\rho\xi}$ and $\overline{\dot{\omega}_{\mathrm{H_{2}}}}$, in order to calculate $\widetilde{c}$, $\widetilde{\phi}$ and $\overline{\dot{\omega}}$ (Eqs (2.3), (2.6) and (2.4)). The filtered quantity of a variable $Q$ from the DNS field is defined as
[…]
where $F$ is the spatial filter. The filtered 3D fields are obtained by successively filtering on each direction $x$, $y$ and $z$ using a One-Dimensional (1D) Gaussian filter. The 1D filter function is written in discrete form as
@@ 3.2 Generalization ability
Figure 7: Scatter plots with 2D histograms: CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ versus ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. Individual values are normalized by the maximum burning rate in the datasets. The points used for the histograms have a progress variable $c$: $0.05\leq c\leq 0.95$. Histogram values below the colour scale are transparent. Gray dashed line indicates $x=y$ (i.e. zero error). Each column corresponds to a global equivalence ratio. Each row corresponds to a set of LES parameters (filtering and downsampling). Data are collected from solutions with two sets of LES parameters that were not included into the training dataset. The CNN approximates the burning rates with good accuracy, demonstrating the ability to generalise to other LES parameters from which it has been trained
@@
Figure 8: Scatter plots and 2D histograms together with planar cuts comparing CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ to ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. See caption of Fig. 5 for a full description of how the histograms are constructed. The two global equivalence ratios $\phi_{g}=0.45$ and $0.55$ were not included in the training dataset. The CNN approximates the burning rates with high accuracy, demonstrating the ability to generalise to other equivalence ratios from which it has been trained
@@ 3.3 Comparison with a filtered tabulated chemistry approach
A filtered tabulated chemistry modeling is implemented to compare CNN results with an established approach for modeling filtered burning rates. In the filtered tabulated chemistry framework, the flame structure in the direction normal to the flame front is assumed identical to the structure of a planar 1D freely propagating premixed laminar flame [11]. For a given filter size, it is then possible to apply the filter operation introduced in Section 2.3.2 (Eqs (2.8) and (2.9)) to a precomputed 1D flame and extract the burning rates directly as
$\overline{\dot{\omega}}=\overline{\dot{\omega}}^{\,+}\left(\widetilde{c}\right)\,\text{,}$ (3.2)
where $\overline{\dot{\omega}}^{\,+}$ is the filtered burning rate of the 1D flame, mapped as a function of the filtered progress variable $\widetilde{c}$ for a given filter size $\sigma$. Flame wrinkling at the subfilter-scale level is lost in LES. A subfilter-scale wrinkling factor $\Xi$ needs to be applied to compensate for the loss in flame surface [30]. The filtered burning rate for filtered tabulated chemistry modeling therefore reads
$\overline{\dot{\omega}}^{FC}=\Xi\,\overline{\dot{\omega}}^{\,+}\left(\widetilde{c}\right)\,\text{.}$ (3.3)
The wrinkling factor $\Xi$ is usually modeled by algebraic expressions. Models based on a fractal description of the flame front [14, 15] are very common. They assume that in a range of physical scales bounded by an inner cutoff $\eta$ and an outer cutoff $L$, the flame front is a fractal surface of dimension $2\leq D_{f}\leq 3$. The wrinkling factor is given by
@@
$L$ corresponds to the size of the largest unresolved wrinkles, which is of the order of the thickness of the LES flame estimated as $\delta_{c,L}^{1}$ (Eq. (2.10)). $\eta$ is the size of the smallest wrinkles and varies from $\eta=L$ (i.e. no unresolved wrinkling, $\Xi=1$) to $\eta=\delta_{c,L}^{0}$ (i.e. filling of the space by a flame surface of fractal dimension $D_{f}$ between cut-off scales $\delta_{c,L}^{0}$ and $L$, $\Xi=\left(L/\delta_{c,L}^{0}\right)^{D_{f}-2}$). The latter situation is reached when turbulence intensity is sufficiently high, which is often the case in practice [40, 39]. We then assume $\eta=\delta_{c,L}^{0}$ in Eq. (3.4), maximizing the flame surface density at the subfilter-scale level. The fractal dimension is estimated as $D_{f}=2.5$ following the work of Charlette et al. [8].
CNN-based modeling significantly outperforms the filtered tabulated chemistry one, especially for the lean cases (e.g. $\phi_{g}=0.35$ in Fig. 9). Due to the extremely high thermodiffusive effects in the very lean cases, the burning rates of the turbulent 3D field can be much higher than those in a freely propagating planar flame at the same global equivalence ratio $\phi_{g}$. For example, for the case $\phi_{g}=0.35$, $\sigma=4$, $\mathrm{DSF}=2$, the maximum burning rate in the 3D field $\mathrm{max}(\overline{\dot{\omega}}^{\ast})$ is about $17$ times greater than that in a laminar planar flame $\mathrm{max}(\overline{\dot{\omega}}^{\,+})$. This is due to local stretching and enrichment effects that are not taken into account by the tabulated chemistry approach we have implemented. The wrinkling factor $\Xi$ only accounts for flame surface density, and new models need to be developed. This $\phi_{g}=0.35$ case is a truly extreme test case to illustrate the strength of the CNN model. Tabulated chemistry modeling does not fail so spectacularly for equivalence ratios close to stoichiometry, where preferential and differential diffusion effects are less pronounced (e.g. $\phi_{g}=0.7$ in Fig. 9). The CNN results are still much more accurate thanks to the consideration of the 3D flame structure via successive convolutions.
@@
Figure 9: Scatter plots of the CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ and the filtered tabulated chemistry burning rate $\overline{\dot{\omega}}^{\mathrm{FC}}$ (Eq. (3.3)) versus ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. Individual values are normalized by the maximum burning rate in the datasets. Gray dashed line indicates $x=y$ (i.e. zero error). Two cases of LES parameter sets are presented for two global equivalence ratios: the lowest equivalence ratio $\phi_{g}=0.35$ for which the filtered tabulated chemistry is significantly flawed due to very strong thermodiffusive effects; the highest equivalence ratio $\phi_{g}=0.7$ for which thermodiffusive effects are less important
@@ 4 Conclusion and future work
The trained CNN model retrieves burning rates with very high accuracy on the 3D full-scale test solutions. It is able to model burning rates in a very complex turbulent flame topology with flame elements detached from the main flame brush. In addition, the CNN approximates burning rates with low errors for filter and downsampling parameters and global equivalence ratios other than those for which it has been trained. The ML-based modeling outperforms a priori a classical filtered tabulated chemistry approach with correction for flame surface density due to subfilter-scale flame wrinkling. This is particularly true for very lean flames, where preferential diffusion effects are major and undermine modeling based on flame surface density.
@@
These results are extremely promising. They demonstrate that it is possible to build an ML-based model of the burning rates of turbulent lean premixed H2-air flames—an extremely challenging task—that can generalise outside the scope for which it has been trained. Generalisation is very important if we are really going to consider implementing this type of model in engineering simulation software. This work paves the way for a new modeling approach to turbulent H2 combustion, which we want to pursue with the scientific community in the very near future.
@@
We can then draw up a roadmap for integration into an LES code. First, the modeling of the diffusive fluxes remains to be addressed. It is usually done through simple gradient assumptions [30]. However, since diffusion plays an important role in the propagation speed of premixed flames [41, 18], a simple closure of this type may not be sufficient. If this is the case, we could consider training the CNN to approximate the diffusive fluxes using several output channels. Another option could be to derive analytical laws to correct the diffusive fluxes so as to retrieve appropriate flame propagation speeds. Second, there are a number of practical integration challenges to be tackled. In a massively parallel HPC code, the fluid domain is decomposed in many partitions distributed over the MPI tasks (typically one MPI per GPU). When inference is performed on each MPI, artefacts may appear at the boundaries of the partitions [37]. To avoid this, the inference must be given its neighboring context to produce overlapping predictions. However, this incurs additional MPI communication costs that must be evaluated on a case by case basis. Another major challenge is related to the voxel grid requirement for the CNN. If the CFD code uses an irregular grid that does not match the voxel/pixel structure, two grids have to be used: one for the fluid solver, one for the CNN inference. This requires the development of efficient methods for interpolating on-the-fly and transferring data between the two grids. The PhyDLL [^1] https://phydll.readthedocs.io [37] open-source library is one effort that could help to address these issues. These questions will be the focus of a future project, now that we have shown that CNN-based modeling of the burning rates is possible with this work.
@@ Acknowledgments
The DNS database will be made available via BLASTNet https://blastnet.github.io/index.html.

### v2

@@ 2.3.2 Filtering and downsampling
A spatial filtering is performed to obtain $\overline{\rho}$, $\overline{\rho Y_{\mathrm{H_{2}}}}$, $\overline{\rho\xi}$ and $\overline{\dot{\omega}_{\mathrm{H_{2}}}}$, in order to calculate $\widetilde{c}$, $\widetilde{\phi}$ and $\overline{\dot{\omega}}$ (Eqs (2.3), (2.5) and (2.1)). The filtered quantity of a variable $\varphi$ from the DNS field is defined by Eq. (1.1). The filtered 3D fields are obtained by successively filtering on each direction $x$, $y$ and $z$ using a 1D Gaussian filter. The 1D filter function is written in discrete form as
@@ 3.2 Generalization ability
Figure 7: Scatter plots with 2D histograms: CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ versus ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. Individual values are normalized by the maximum burning rate in the datasets. The points used for the histograms have a progress variable $c$: $0.05\leq c\leq 0.95$. Histogram values below the colour scale are transparent. Gray dashed line indicates $x=y$ (i.e. zero error). Each column corresponds to a global equivalence ratio. Each row corresponds to a set of LES parameters (filtering and downsampling). Data are collected from solutions with two sets of LES parameters that were not included into the training dataset. The CNN approximates the burning rates with good accuracy, demonstrating the ability to generalize to other LES parameters from which it has been trained
@@
Figure 8: Scatter plots and 2D histograms together with planar cuts comparing CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ to ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. See caption of Fig. 5 for a full description of how the histograms are constructed. The two global equivalence ratios $\phi_{g}=0.45$ and $0.55$ were not included in the training dataset. The CNN approximates the burning rates with high accuracy, demonstrating the ability to generalize to other equivalence ratios from which it has been trained
@@ 3.3 Comparison with a filtered tabulated chemistry approach
A filtered tabulated chemistry modeling is implemented to compare CNN results with an established approach for modeling filtered burning rates. In the filtered tabulated chemistry framework, the flame structure in the direction normal to the flame front is assumed identical to the structure of a planar 1D freely propagating premixed laminar flame [20]. For a given filter size, it is then possible to apply the filter operation introduced in Section 2.3.2 (Eqs (1.1) and (2.8)) to a precomputed 1D flame at a given global equivalence ratio $\phi_{g}$ and extract the burning rates directly as
$\overline{\dot{\omega}}\approx\overline{\dot{\omega}}^{\,+}\left(\widetilde{c},\phi_{g}\right)\,\text{,}$ (3.2)
where $\overline{\dot{\omega}}^{\,+}$ is the filtered burning rate of the 1D flame, mapped as a function of the filtered progress variable $\widetilde{c}$ for a given filter size $\sigma$. To account for the local fluctuation of the equivalence ratio in the domain, it is possible to retrieve the reaction rate in a 1D flame computed with the local equivalence ratio $\phi$ used to define the fresh gas composition. In this case, we write
$\overline{\dot{\omega}}\approx\overline{\dot{\omega}}^{\,+}\left(\widetilde{c},\widetilde{\phi}\right)\,\text{.}$ (3.3)
Flame wrinkling at the subfilter-scale level is lost in LES. A subfilter-scale wrinkling factor $\Xi$ needs to be applied to compensate for the loss in flame surface [42]. The filtered burning rate for filtered tabulated chemistry modeling therefore reads
$\overline{\dot{\omega}}^{F}=\Xi\,\overline{\dot{\omega}}^{\,+}\left(\widetilde{c},\phi_{g}\right)\;\text{or}\;\overline{\dot{\omega}}^{FC}=\Xi\,\overline{\dot{\omega}}^{\,+}\left(\widetilde{c},\widetilde{\phi}\right)\;\text{,}$ (3.4)
depending on whether the global or local equivalence ratio is used to define the fresh gas composition in the corresponding 1D flame. The wrinkling factor $\Xi$ is usually modeled by algebraic expressions. Models based on a fractal description of the flame front [24, 25] are very common. They assume that in a range of physical scales bounded by an inner cutoff $\eta$ and an outer cutoff $L$, the flame front is a fractal surface of dimension $2\leq D_{f}\leq 3$. The wrinkling factor is given by
@@
$L$ corresponds to the size of the largest unresolved wrinkles, which is of the order of the thickness of the LES flame estimated as $\delta_{c,L}^{1}$ (Eq. (2.9)). $\eta$ is the size of the smallest wrinkles and varies from $\eta=L$ (i.e. no unresolved wrinkling, $\Xi=1$) to $\eta=\delta_{c,L}^{0}$ (i.e. filling of the space by a flame surface of fractal dimension $D_{f}$ between cut-off scales $\delta_{c,L}^{0}$ and $L$, $\Xi=\left(L/\delta_{c,L}^{0}\right)^{D_{f}-2}$). The latter situation is reached when turbulence intensity is sufficiently high, which is often the case in practice [56, 55]. We then assume $\eta=\delta_{c,L}^{0}$ in Eq. (3.5), maximizing the flame surface density at the subfilter-scale level. The fractal dimension is estimated as $D_{f}=2.5$ following the work of Charlette et al. [13]. The set of 1D flames are computed using the San Diego mechanism [45] for the chemical kinetics (same as the training database for the CNN) and a mixture-averaged transport model for the diffusivities.
CNN-based modeling significantly outperforms the filtered tabulated chemistry one, especially for the lean cases (e.g. $\phi_{g}=0.35$ in Fig. 9). Due to the extremely high thermodiffusive effects in the very lean cases, the burning rates of the turbulent 3D field can be much higher than those in a freely propagating planar flame at the same equivalence ratio. For example, for the case $\phi_{g}=0.35$, $\sigma=4$, $\mathrm{DSF}=2$, the maximum burning rate in the 3D field $\mathrm{max}(\overline{\dot{\omega}}^{\ast})$ is about $17$ times greater than that in a laminar planar flame $\mathrm{max}(\overline{\dot{\omega}}^{\,+})$. This is due to the fact that flame strain is not considered in the tabulation method applied, and that the subfilter wrinkling factor $\Xi$ only takes into account turbulence-induced wrinkling, ignoring that induced by thermodiffusive instabilities and their interactions with turbulence [7]. This $\phi_{g}=0.35$ case is a truly extreme test case to illustrate the strength of the CNN model. Tabulated chemistry modeling does not fail so spectacularly for equivalence ratios close to stoichiometry, where preferential and differential diffusion effects are less pronounced [6] (e.g. $\phi_{g}=0.7$ in Fig. 9). The CNN results are still much more accurate thanks to the consideration of the 3D flame structure via successive convolutions. To further demonstrate the CNN ability to learn from flame topology, training has been performed without information on the equivalence ratio for the case $\phi_{g}=0.35$. This study is given as supplementary material. It is found that burning rate approximation is still very accurate, proving our initial claim that a CNN is capable of recognizing complex flame topology and inferring associated burning rates by learning spatial hierarchies of features.
@@
Figure 9: Scatter plots of the CNN-modeled burning rate $\overline{\dot{\omega}}^{\mathrm{NN}}$ and the filtered tabulated chemistry burning rate $\overline{\dot{\omega}}^{\mathrm{F}}$ or $\overline{\dot{\omega}}^{\mathrm{FC}}$ (Eq. (3.4)) versus ground-truth filtered burning rate $\overline{\dot{\omega}}^{\ast}$. Individual values are normalized by the maximum burning rate in the datasets. Gray dashed line indicates $x=y$ (i.e. zero error). Two cases of LES parameter sets are presented for two global equivalence ratios: the lowest equivalence ratio $\phi_{g}=0.35$ for which the filtered tabulated chemistry is significantly flawed due to very strong thermodiffusive effects; the highest equivalence ratio $\phi_{g}=0.7$ for which thermodiffusive effects are less important
@@ 4 Conclusion and future work
The trained CNN model retrieves burning rates with very high accuracy on the 3D full-scale test solutions. It is able to model burning rates in a very complex turbulent flame topology with flame elements detached from the main flame brush. In addition, the CNN approximates burning rates with low errors for filter and downsampling parameters and global equivalence ratios other than those for which it has been trained. The ML-based modeling outperforms a priori a classical filtered tabulated chemistry approach with correction for flame surface density due to subfilter-scale flame wrinkling. This is particularly true for very lean flames, where differential diffusion effects are major and undermine modeling based on flame surface density.
@@
These results are extremely promising. They demonstrate that it is possible to build an ML-based model of the burning rates of turbulent lean premixed H2-air flames—an extremely challenging task—that can generalize outside the scope for which it has been trained. Generalization is very important if we are really going to consider implementing this type of model in engineering simulation software. This work paves the way for a new modeling approach to turbulent H2 combustion, which we want to pursue with the scientific community in the very near future.
We can then draw up a roadmap for integration into an LES code. First, the modeling of the subfilter-scale fluxes in Eq. (1.2) remains to be addressed. It could be done through gradient transport models with a turbulent viscosity hypothesis [11]. If this closure fails to retrieve correct flame propagation speeds, we could consider training the CNN to approximate the fluxes using several output channels. Second, the evolution of the entire chemical system must be described. This can be done by training the CNN on the source term of a progress variable, transported instead of the individual species that make up the reactive mixture. This is done, for example, in Refs [29, 30] with a tabulation method for the closure of the transport equation of the progress variable. However, the transport of the mixture fraction must be modeled to account for equivalence ratio fluctuations. We could also train the model again on a database generated with a single global chemical reaction [35, 46]. This has the advantage of being able to transport physical species instead of a progress variable, accounting for preferential diffusion through the multi-species Navier-Stokes equations directly.
@@
Second, there are a number of practical integration challenges to be tackled. In a massively parallel HPC code, the fluid domain is decomposed in many partitions distributed over the MPI tasks (typically one MPI per GPU). When inference is performed on each MPI, artefacts may appear at the boundaries of the partitions [50]. To avoid this, the inference must be given its neighboring context to produce overlapping predictions. However, this incurs additional MPI communication costs that must be evaluated on a case by case basis. Another major challenge is related to the voxel grid requirement for the CNN. If the CFD code uses an irregular grid that does not match the voxel/pixel structure, two grids have to be used: one for the fluid solver, one for the CNN inference. This requires the development of efficient methods for interpolating on-the-fly and transferring data between the two grids. The PhyDLL [^1] https://phydll.readthedocs.io [50] open-source library is one effort that could help to address these issues. These questions will be the focus of a future project, now that we have shown that CNN-based modeling of the burning rates is possible with this work.
@@ Acknowledgments
Code for training and inference is available via GitLab at https://gitlab.com/male.quentin/cnn_h2flame. A set of solutions from DNS is distributed via Kaggle (https://www.kaggle.com) with the following identifiers:
  • malquentin/premixed-flame-slot-burner-dns-h2air-phi035
  • malquentin/premixed-flame-slot-burner-dns-h2air-phi04
  • malquentin/premixed-flame-slot-burner-dns-h2air-phi05
  • malquentin/premixed-flame-slot-burner-dns-h2air-phi06
  • malquentin/premixed-flame-slot-burner-dns-h2air-phi07
The DNS database will be integrated into BLASTNet (https://blastnet.github.io/index.html).